OpenStudy (anonymous):

[6.03] Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 4x + 3y = -1 3x + y = 3

OpenStudy (anonymous):

It would be much easier to use the elimination method in this case. The reason is because if you use the substitution method you would have to isolate a variable and that would make any of these two equations way too complicated. Multiply the bottom equation by -3 to get -3y. Then you add the equations together and solve for x. Then once you have x, plug in x into any of the equations and you'll get Y. 4x + 3y = 1 -9x - 3xy = -9 -5x=-8 x= 1.6 Plug in 1.6 for x: 3(1.6) + y = 3 4.8 + y = 3 y= 3-4.8 y= - 1.8

OpenStudy (anonymous):

is that all your question ?

OpenStudy (anonymous):

Yes..or..No ?

OpenStudy (anonymous):

i got more but can i email you or somethign because i appericate your help my teacher dont help

OpenStudy (anonymous):

so you can ask me for help when you need it ? Because i'm about to leave in a minute.